OpenStudy (anonymous):

I WILL FAN AND MAEDEL !!!!!!!!!!! 26 gallons of gas cost 80.08. what is the unit rate for 1 gallon of gas. A. 2.98 per gallon B. 3.08 per gallon C. 4.16 per gallon D. 308.00 per gallon

OpenStudy (anonymous):

Just divide the cost by the gallons. (: 80.08/26 = 3.08 So, B. 3.08 per gallon

OpenStudy (anonymous):

@BunnyBree is right

OpenStudy (dmndlife24):

Take the total cost and divide it by the number of gallons you got which was 26 26 / 80.08 = 3.08 So B.

OpenStudy (dmndlife24):

* 80.08 / 26 = 3.08

OpenStudy (anonymous):

THANKS BUNNYBREE

OpenStudy (anonymous):

You're welcome! (: